How much of a 50% antifreeze solution must a mechanic mix with a 70% antifreeze solution if 20 gallons of a 65% antifreeze solution are needed?

Answer by mananth(16946)

Percent ---------------- quantity
Antifreeze I 50.00% ---------------- x gallons
Antifreeze II 70.00% ------ 20 - x gallons
Mixture 65.00% ---------------- 20
Total 20 gallons
50.00% x + 70.00% ( 20 - x ) = 65.00% * 20
50 x + 70 ( 20 - x ) = 1300
50 x + 1400 - 70 x = 1300
50 x - 70 x = 1300 - -1400
-20 x = -100
/ -20
x = 5 gallons 50.00% Antifreeze I
15 gallons 70.00% Antifreeze II

m.ananth@hotmail.ca